# Symmetry-Constrained Forecasting of Periodically Correlated Energy Processes

### v1

@@ 3 Modeling Results and Analysis
Two complementary datasets are employed to assess the proposed operators across synthetic and measured conditions. The first consists of 100 controlled cyclostationary series generated through a hybrid $\mathtt{bootstrap}$$\mathtt{Monte\text{-}Carlo}$ scheme, enabling modulation of periodicity, noise, and correlation structure. The second dataset corresponds to an empirical use case in renewable energy forecasting, solar irradiation, based on 30-min observations from 68 $\mathtt{SIAR}$ meteorological stations in Spain [7] (Figure 2, data available in https://observatorioregadio.gob.es/fr/outils/siar/). Spatial consistency is ensured by ordinary kriging, following the empirical variogram of each subregion. Forecast horizons span 30 min to 6 h in half hour increments. An ad hoc quality control following [13] guarantees data integrity. Cyclostationary statistics and learning were estimated using two full years of data. Less than 5% of the observations were missing or rejected during quality control procedures and were excluded from the analysis. The final year serves as out-of-sample validation under an incast configuration, reconstructing past states from subsequent observations to quantify retrospective forecast skill. All methodological steps are fully described through analytical formulations and pseudocode, enabling independent reproduction of the experiments in any programming environment. The $\mathtt{SIAR}$ observational data used for validation are publicly accessible via the official network portal; however, redistribution is subject to the data provider’s policy.
@@ 3.1 Generation of Synthetic Cyclostationary Time Series
The Monte Carlo analysis is based on 100 independent cyclostationary time series. Additional experiments conducted with 500 replications confirmed the convergence and stability of the results. These additional tests did not alter the conclusions and are therefore not shown for brevity. Algorithm 3.1 generates synthetic periodically stationary series $I(t)$, $t=1,\ldots,N$, by combining a deterministic periodic component $\tilde{X}_{t}$ with correlated noise filtered through a low pass smoother. The resulting signals reproduce periodic patterns with controlled stochastic variability, emulating typical cyclostationary behavior. Four statistical indicators characterize each realization: the Coefficient of Variation ($\mathtt{CV}$) for relative variability, the Mean Absolute Return ($\mathtt{MAR}$) for temporal volatility, the Root Mean Square Error ($\mathtt{RMSE}$) for deviation from the periodic trend, and the lag-1 autocorrelation $\rho(1)$ for short term dependence. This design ensures rigorous control of both deterministic and stochastic properties within the generated ensemble.
@@
4:  Bound noise: $\mathtt{WN}(t)\leftarrow\min(\max(\text{WN}(t),0.2),1.1)$ Ensure noise is between 0.2 and $1\triangleright\,\text${Ensure noise is between 0.2 and 1.1}
5:  Apply low-pass filter: $\mathtt{LPF}(\mathtt{WN}(t))\leftarrow\frac{1}{L}\sum_{k=0}^{L-1}\mathtt{WN}(t-k),\quad\text{if }t\geq L$ $\triangleright\,\text{$L$-size window Smoothing}$
@@
[…]
@@ 3.2 Benchmark of Reference Models with Synthetic Data
Among the forecasting operators in Table 1, three reference models were retained for benchmarking: Holt–Winters ($\mathtt{HW}$), the Theta method, and the Periodic Autoregressive model ($\mathtt{PAR}$). These methods are classical analytical frameworks able to represent trends, seasonality, and cyclostationary dependencies while remaining computationally efficient. In contrast, the persistence based operators proposed in this work are parameter free and serve as analytical baselines for evaluating the potential gains achievable with model driven approaches. For $\mathtt{HW}$ and Theta, parameters were optimized via $\mathtt{MATLAB}$’s fmincon Interior Point solver; ten independent runs were averaged to reduce local minima bias. The $\mathtt{PAR}$ model was implemented with order $p=T$, exploiting the intrinsic periodicity of non seasonally adjusted signals. A Monte Carlo protocol of 100 realizations was applied across forecast horizons $n\Delta t\in[1,9]$. Each series used randomized parameters $T\sim\mathcal{U}(10,50)$ and $L\sim\mathcal{U}(1,10)$ (see Algorithm 3.1), to ensure statistical diversity and robustness. Figure 4 shows the relation between normalized RMSE and signal variability, with trends fitted by third order polynomials. As the horizon increases, $\mathtt{nRMSE}$ sensitivity to variability weakens, reflecting the diminishing relevance of short term correlation beyond $\sim$6 h. Figure 5 summarizes the distribution of $\mathtt{nRMSE}$ across models and horizons, confirming the expected performance hierarchy among reference and persistence based operators.
@@
These conclusions remain conditioned by the synthetic experiments, where trend and periodicity are explicitly defined. Persistence based models demonstrate strong efficiency under such controlled conditions, whereas statistical methods such as $\mathtt{Theta}$, $\mathtt{PAR}$, and $\mathtt{HW}$ show limited adaptability to cyclic trends. Subsequent analyses therefore focus on a restricted subset of operators: $\tilde{\mathtt{P}}^{\circlearrowleft}_{\mathtt{BLEND}}$ and $\mathtt{P}^{\circlearrowleft}_{\mathtt{BLEND}}$, which exhibited the most consistent performance; $\mathtt{P}_{\mathtt{CLIPER}}$, retained for its analytical simplicity; and $\mathtt{P_{S}}$ (Smart Persistence), the most widely used naive predictor in operational solar forecasting. For completeness, the canonical persistence models $\mathtt{P}$ and $\mathtt{P}^{\circlearrowleft}$ are also included, since the latter performs notably well at extended horizons. Two lightweight learning based baselines are incorporated for comparison: (i) a linear AutoRegressive model ($\mathtt{AR}$, [17]) applied to stationary transformed data (e.g., clear sky ratio [23]) and (ii) an Extreme Learning Machine ($\mathtt{ELM}$, [34]) trained directly on raw irradiance without normalization. All other complex or hybrid models are intentionally excluded to maintain focus on physically interpretable and analytically tractable formulations. Finally, convexity and boundedness effects identified in Lemmas 35 (App. B & C) further justify the preference for the simplified operator $\tilde{\mathtt{P}}^{\circlearrowleft}_{\mathtt{BLEND}}$, which ensures stability and interpretability without compromising analytical coherence.
@@ Appendix A Complexity of an Algorithm
The complexity of an algorithm measures the time or memory space required as a function of the input size $n$. The $\mathcal{O}(.)$ notation is used to express this complexity by estimating how quickly the execution time grows as $n$ increases. Considering an algorithm with $\mathcal{O}(n)$, its execution time increases proportionally to the input size. For example, in an Autoregressive ($\mathtt{AR}$) model of order $\mathtt{P}$, the next value $I(t+1)$ is forecasted based on a linear combination of the previous $p$ values:
@@
where $\phi_{i}$ are the coefficients and $\epsilon(t)$ is a white noise term. The training complexity of the AR model is $\mathcal{O}(n\mathtt{P}^{2})$ (inversion of $\mathbf{X}^{\top}\mathbf{X}$ in the least squares optimization process), as fitting the model involves solving a system of equations derived from $n$ observations, requiring the inversion of a $p\times p$ matrix. The forecasting complexity is $\mathcal{O}(p)$, as it depends only on the $p$ most recent observations.
@@ Appendix C Derivation for the BLEND Persistence Operator under Cyclostationary and Stationary Assumptions
$\mathtt{MSE}=$
$\mathbb{E}\left[\left(I(t+n\Delta t)-(1-\lambda^{\circlearrowleft}_{\mathtt{BLEND}})I(t-T+n\Delta t)-\lambda^{\circlearrowleft}_{\mathtt{BLEND}}I(t)\right)^{2}\right].$
@@ Appendix D Properties of BLEND Operator Family
Substituting $I(t)=c$ into the definition:

### v3

@@ 3 Modeling Results and Analysis
Two complementary datasets are employed to assess the proposed operators across synthetic and measured conditions. The first consists of 100 controlled cyclostationary series generated through a hybrid $\mathtt{bootstrap}$$\mathtt{Monte\text{-}Carlo}$ scheme, enabling modulation of periodicity, noise, and correlation structure. The second dataset corresponds to an empirical use case in renewable energy forecasting, solar irradiation, based on 30-min observations from 68 $\mathtt{SIAR}$ meteorological stations in Spain [10] (Figure 2, data available in https://observatorioregadio.gob.es/fr/outils/siar/). Spatial consistency is ensured by ordinary kriging, following the empirical variogram of each subregion. Forecast horizons span 30 min to 6 h in half hour increments. An ad hoc quality control following [16] guarantees data integrity. Cyclostationary statistics and learning were estimated using two full years of data. Less than 5% of the observations were missing or rejected during quality control procedures and were excluded from the analysis. The final year serves as out-of-sample validation under an incast configuration, reconstructing past states from subsequent observations to quantify retrospective forecast skill. All methodological steps are fully described through analytical formulations and pseudocode, enabling independent reproduction of the experiments in any programming environment. The $\mathtt{SIAR}$ observational data used for validation are publicly accessible via the official network portal; however, redistribution is subject to the data provider policy. All MATLAB source code implementing the operators described in this section is openly available at https://github.com/cyrilvoyant/cyclostationary-forecasting-matlab (DOI: https://doi.org/10.5281/zenodo.18812334), enabling full reproducibility of the presented results.
@@ 3.1 Generation of Synthetic Cyclostationary Time Series
The Monte Carlo analysis is based on 100 independent cyclostationary time series. Additional experiments conducted with 500 replications confirmed the convergence and stability of the results. These additional tests did not alter the conclusions and are therefore not shown for brevity. Algorithm 3.1 generates synthetic periodically stationary series $I(t)$, $t=1,\ldots,N$, by combining a deterministic periodic component $\tilde{X}_{t}$ with correlated noise filtered through a low pass smoother.
@@
4:  Bound noise: $\mathtt{WN}(t)\leftarrow\min(\max(\mathtt{WN}(t),0.2),1.1)$ Ensure noise is between 0.2 and $1\triangleright\,\text${Ensure noise is between 0.2 and 1.1}
5:  Low-pass filter: $\mathtt{LPF}(\mathtt{WN}(t))\leftarrow\frac{1}{L}\sum_{k=0}^{L-1}\mathtt{WN}(t-k),\quad\text{if }t\geq L$ $\triangleright\,\text{$L$-size window Smoothing}$
@@
The resulting signals reproduce periodic patterns with controlled stochastic variability, emulating typical cyclostationary behavior. Four statistical indicators characterize each realization: the Coefficient of Variation ($\mathtt{CV}$) for relative variability, the Mean Absolute Return ($\mathtt{MAR}$) for temporal volatility, the Root Mean Square Error ($\mathtt{RMSE}$) for deviation from the periodic trend, and the lag-1 autocorrelation $\rho(1)$ for short term dependence. This design ensures rigorous control of both deterministic and stochastic properties within the generated ensemble. Figure 3 illustrates cyclic parameters and forecasting performance for a noisy periodic signal with period $T=24$. The left panel shows the cyclic mean, standard deviation, and correlation coefficient (undefined for all-zero phase bins), each in raw and smoothed form to highlight the periodic structure. The right panel compares representative forecasting operators: $\mathtt{P}$, $\mathtt{P}^{\circlearrowleft}$, $\mathtt{P_{S}}$, $\mathtt{P}_{\mathtt{CLIPER}}$, $\mathtt{P}^{\circlearrowleft}_{\mathtt{CLIPER}}$, $\tilde{\mathtt{P}}^{\circlearrowleft}_{\mathtt{CLIPER}}$, $\mathtt{P}_{\mathtt{BLEND}}$, $\mathtt{P}^{\circlearrowleft}_{\mathtt{BLEND}}$, and $\tilde{\mathtt{P}}^{\circlearrowleft}_{\mathtt{BLEND}}$. These results demonstrate the operators’ ability to capture both short term dependencies and periodic structures. Cyclic parameters summarize the statistical organization of the process, while the comparative forecasts highlight differences in predictive skill and temporal symmetry preservation.
@@ 3.2 Benchmark of Reference Models with Synthetic Data
Among the forecasting operators in Table 1, three reference models were retained for benchmarking: Holt–Winters ($\mathtt{HW}$), the Theta method, and the Periodic Autoregressive model ($\mathtt{PAR}$). These methods are classical analytical frameworks able to represent trends, seasonality, and cyclostationary dependencies while remaining computationally efficient. In contrast, the persistence based operators proposed in this work are parameter free and serve as analytical baselines for evaluating the potential gains achievable with model driven approaches. For $\mathtt{HW}$ and Theta, parameters were optimized via $\mathtt{MATLAB}$ fmincon Interior Point solver; ten independent runs were averaged to reduce local minima bias. The $\mathtt{PAR}$ model was implemented with order $p=T$, exploiting the intrinsic periodicity of non seasonally adjusted signals. A Monte Carlo protocol of 100 realizations was applied across forecast horizons $n\Delta t\in[1,9]$. Each series used randomized parameters $T\sim\mathcal{U}(10,50)$ and $L\sim\mathcal{U}(1,10)$ (see Algorithm 3.1), to ensure statistical diversity and robustness. Figure 4 shows the relation between normalized RMSE and signal variability, with trends fitted by third order polynomials.
@@
As the horizon increases, $\mathtt{nRMSE}$ sensitivity to variability weakens, reflecting the diminishing relevance of short term correlation beyond $\sim$6 h. Figure 5 summarizes the distribution of $\mathtt{nRMSE}$ across models and horizons, confirming the expected performance hierarchy among reference and persistence based operators.
@@
These conclusions remain conditioned by the synthetic experiments, where trend and periodicity are explicitly defined. Persistence based models demonstrate strong efficiency under such controlled conditions, whereas statistical methods such as $\mathtt{Theta}$, $\mathtt{PAR}$, and $\mathtt{HW}$ show limited adaptability to cyclic trends. Subsequent analyses therefore focus on a restricted subset of operators: $\tilde{\mathtt{P}}^{\circlearrowleft}_{\mathtt{BLEND}}$ and $\mathtt{P}^{\circlearrowleft}_{\mathtt{BLEND}}$, which exhibited the most consistent performance; $\mathtt{P}_{\mathtt{CLIPER}}$, retained for its analytical simplicity; and $\mathtt{P_{S}}$ (Smart Persistence), the most widely used naive predictor in operational solar forecasting. For completeness, the canonical persistence models $\mathtt{P}$ and $\mathtt{P}^{\circlearrowleft}$ are also included, since the latter performs notably well at extended horizons. Two lightweight learning based baselines are incorporated for comparison: (i) a linear AutoRegressive model ($\mathtt{AR}$, [20]) applied to stationary transformed data (e.g., clear sky ratio [27]) and (ii) an Extreme Learning Machine ($\mathtt{ELM}$, [37]) trained directly on raw irradiance without normalization. All other complex or hybrid models are intentionally excluded to maintain focus on physically interpretable and analytically tractable formulations. Finally, convexity and boundedness effects identified in Lemmas 35 (B & C) further justify the preference for the simplified operator $\tilde{\mathtt{P}}^{\circlearrowleft}_{\mathtt{BLEND}}$, which ensures stability and interpretability without compromising analytical coherence.
@@ Appendix A Complexity of an Algorithm
The complexity of an algorithm measures the time or memory space required as a function of the input size $n$. The $\mathcal{O}(.)$ notation is used to express this complexity by estimating how quickly the execution time grows as $n$ increases. Considering an algorithm with $\mathcal{O}(n)$, its execution time increases proportionally to the input size. For example, in an Autoregressive ($\mathtt{AR}$) model of order $p$, the next value $I(t+1)$ is forecasted based on a linear combination of the previous $p$ values:
@@
where $\phi_{i}$ are the coefficients and $\epsilon(t)$ is a white noise term. The training complexity ($n\gg p$) of the $\mathtt{AR}$ model is $\mathcal{O}(n\mathtt{P}^{2})$ (inversion of $\mathbf{X}^{\top}\mathbf{X}$ in the least squares optimization process), as fitting the model involves solving a system of equations derived from $n$ observations, requiring the inversion of a $p\times p$ matrix. The forecasting complexity is $\mathcal{O}(p)$, as it depends only on the $p$ most recent observations.
A.1 Recursive state-space models (HW, ETS, Theta)
These models rely on recursive updates of a small number of state variables (level $L(t)$, trend $b(t)$, seasonal component $S(t)$). Each update involves a constant number of operations. Training requires a single pass through the data $\mathcal{C}_{T}=\mathcal{O}(n).$ Forecasting over $n$ steps is performed via recursive propagation, $\mathcal{C}_{F}=\mathcal{O}(n).$ For the Theta method, $k$ transformed series are processed, so $\mathcal{C}_{T}=\mathcal{O}(nk)\text{and}\,\mathcal{C}_{F}=\mathcal{O}(n).$
A.2 Periodic autoregressive models (PAR)
The $\mathtt{PAR}$ model involves phase-dependent coefficients estimated separately for each phase. This leads to a training complexity dominated by $\mathcal{O}(np^{2})$ when $n/T\gg p$ (see $\mathtt{AR}$ exemple at the begining of the section). Forecasting requires a linear combination of the $p$ most recent values $\mathcal{C}_{F}=\mathcal{O}(p).$
A.3 Regression-based models (Prophet)
Prophet is based on additive regression combining trend and seasonal components, typically represented using $k$ Fourier terms. Training involves the construction of a design matrix of size $n\times k$ and parameter estimation through optimization procedures, leading to $\mathcal{C}_{T}=\mathcal{O}(nk).$ Forecasting consists in evaluating the fitted model over the prediction horizon. Each prediction requires evaluating the trend and seasonal components, resulting in $\mathcal{C}_{F}=\mathcal{O}(nk),$ which reduces to $\mathcal{O}(n)$ when $k$ is fixed.
A.4 State-space models with seasonal differencing (SARIMA)
$\mathtt{SARIMA}$ models involve autoregressive and moving average components with seasonal differencing. Training relies on parameter estimation procedures combining iterative optimization and matrix operations. The dominant computational cost arises from processing $n$ observations with $p$-dimensional structures, leading to $\mathcal{C}_{T}=\mathcal{O}(np^{2}),$ ($n\gg p$). Forecasting is driven by the autoregressive structure and requires evaluating $p$ past terms $\mathcal{C}_{F}=\mathcal{O}(p).$
A.5 Models with implementation-dependent complexity (STL, TBATS, BEATS)
For $\mathtt{STL}$, $\mathtt{TBATS}$, and $\mathtt{N\text{-}BEATS}$, both training and forecasting complexities depend strongly on implementation choices (e.g., smoothing parameters, optimization procedures, or network architecture). As a result, no fixed computational complexity is reported for these models.
@@ Appendix C Derivation for the BLEND Persistence Operator under Cyclostationary and Stationary Assumptions
$\mathtt{MSE}=\mathbb{E}\left[\left(I(t+n\Delta t)-(1-\lambda^{\circlearrowleft}_{\mathtt{BLEND}})I(t-T+n\Delta t)-\lambda^{\circlearrowleft}_{\mathtt{BLEND}}I(t)\right)^{2}\right].$
@@ Appendix D Properties of BLEND Operator Family
Substituting $I(t)=c$ into the definition 15, it comes: